OpenStudy (anonymous):

1. Find the general solution (or as close as you can come to it) for the following differential equations, using separation of variables. (a) dy/dt = 2y − 1 (b) dy/dt = t^2y^3

OpenStudy (dumbcow):

\[\frac{dy}{dt} -2y = -1\] find integrating factor \[e^{\int\limits_{}^{}-2} = e^{-2t}\] \[\frac{dy}{dt}e^{-2t} -2e^{-2t}y = -e^{-2t}\] \[(e^{-2t}y)\frac{d}{dt} = -e^{-2t}\] integrate both sides with respect to t \[e^{-2t}y = \frac{1}{2}e^{-2t} + C\] \[y = \frac{1}{2} +Ce^{2t}\]

OpenStudy (dumbcow):

\[\frac{dy}{dt} = t^{2}y^{3}\] separate variables \[\frac{dy}{y^{3}} = t^{2} dt\] integrate both sides \[-\frac{1}{2y^{2}} = \frac{t^{3}}{3}+C\] \[2y^{2} = \frac{3}{C-t^{3}}\] \[y = \sqrt{\frac{3}{C-2t^{3}}}\]